OpenStudy (summersnow8):

Help anyone? The Power You Produce: Estimate the power you produce in running up a flight of stairs. Give your answer in horsepower. the power we expend during flight is W= 2.3MJ

OpenStudy (festinger):

1hp=746W. Without the knowledge of how high the stairs is, I can only give you at outline. Take 2.3*10^6J and divide it by the time esimate, then take this new value and divide by 746W to get the answer.

OpenStudy (anonymous):

2.3 Mj is work done during running up stairs... time will give u power.. than u can convert in horsepower by dividing with 746

OpenStudy (anonymous):

w/t=power

OpenStudy (theeric):

Pretty much what they said. You'll have to explain your estimation, right? Your approach has to be scientific so that your answer more probably correct. So, what they said: Power is \(\Large P=\frac{W}{t}\) Where \(P\) is power, \(W\) is the work (given), and \(t\) is the time it took to get up the stairs. So, that's the first part of the science. Then you'll go into how you find the time. Time yourself going up the stairs. If you want, time yourself multiple times and take the average. If you time yourself multiple times, you should take a break between times so that being tired out isn't a factor. So you'll explain how you got your time. Then you'll use all that science to arrive at your estimate, where \(W\) is given and you found \(t\). Then you'll have your estimate.

OpenStudy (theeric):

And since \(W\) is in MJ, or \(10^6 J\), and you'll measure \(t\) in seconds, you'll have \(P\) in units of \(10^6\ [J/s]=10^6\ [\text{watts}]\). You can convert, using information Festinger supplied:\[1\ [hp]=746\ [\text{watts}]\]\[\qquad\Downarrow\]\[\frac{1\ [hp]}{746\ [\text{watts}]}=1\]

OpenStudy (summersnow8):

ok I think i get it

OpenStudy (theeric):

Cool! Any questions?